OpenStudy (anonymous):

Alan finished his homework in 1 1/2 hours. It took jimmy 3/4 of an hour longer than alan to finish his homework. How long did it take Jimmy to finish his homework?

hero (hero):

Add 1 + 1/2 + 3/4

OpenStudy (anonymous):

\[2\frac{ 1 }{ 4 } \] hours or 135 minutes

hero (hero):

1 + 1/2 + 3/4 = 1 + 2/4 + 3/4 = 1 + 5/4 = 1 + 4/4 + 1/4 = 1 + 1 + 1/4 = 2 + 1/4 (in hours)